OpenStudy (anonymous):

Solve: 6(x + 4) = 8(x − 2)

OpenStudy (anonymous):

6(x+4)=8(x-2) = 6x + 24 = 8x - 16 8x - 6x = 24 + 16 2x = 40 x = 20

OpenStudy (anonymous):

6(x+4) = 8(x-2) 6x+24 = 8x-16 -6x -6x 24=2x-16 +16 +16 40=2x X=20